# F-95000 Cergy-Pontoise, France

### v1

@@ 3.2 True, Consistent and Conflicting Tuples in a Star Schema
Figure 2: The star-table $T$ and the m-table $m\_Chase(T)$ of Example 3
In the context of Example 3, the star-table $T$ and its associated m-table $m\_Chase(T)$ are shown in Figure 2. In particular, it can be seen that, in $m\_Chase(T)$, the last m-tuple satisfies item 3.a of Proposition 2 while the first three m-tuples satisfy item 3.b. As will be seen later, these three m-tuples are the ones to be relevant for analytic queries.
@@
Using Proposition 1, the set ${\sf Confl}_{\min}(\mathcal{T})$ is characterized as follows: $t$ is in ${\sf Confl}_{\min}(\mathcal{T})$ if and only if one of the following two statements holds:
  • there exist $\sigma$ in $m\_Chase(T)$, $i_{0}$ in $\{1,\ldots,n\}$ and $A$ in $sch^{*}(D_{i_{0}})\cap sch(\sigma)\cap sch(t)$ such that $t=k_{i_{0}}a$ is in ${\sf tuples}(\sigma(K_{i_{0}}A))$ and $|{\sf tuples}(\sigma(A))|>1$
  • there exists $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma)\cap sch(t)$, there exists $M_{i}$ in $\mathbf{M}\cap sch(\sigma)\cap sch(t)$ such that $t=km_{i}$ is in ${\sf tuples}(\sigma(\mathbf{K}M_{i}))$ and $|{\sf tuples}(\sigma(M_{i}))|>1$.
@@
Applying Proposition 1 in the context of Example 3, for which Figure 2 displays the star-table $T$ and the m-table $m\_Chase(T)$, the sets ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$ are as follows:
  • ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of the tuples in ${\sf tuples}(\sigma)$ for every $\sigma$ in $m\_Chase(T)$. Thus, ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of:
    -- $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    -- $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
    -- $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
    -- $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
    -- $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    -- $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$.
    -- $k_{1}k_{2}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    -- $k_{1}k^{\prime}_{2}a_{2}b_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
    -- $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
    -- $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$. $\Box$
@@ 4.1 Definition of Repairs
As explained in LauS25 , contrary to most approaches to consistent query answering in the literature, it is not appropriate to define a repair of $T$ as a maximal subset of ${\sf True}(\mathcal{T})$ satisfying $FD$. This is so because it is intuitively justified to define a repair $R$ of $T$ so as ${\sf True}(\mathcal{R})$ contains ${\sf Cons}(\mathcal{T})$. Unfortunately as shown earlier in Example 2, it may happen that a maximal consistent subset S of ${\sf True}(\mathcal{T})$ does not contain all consistent tuples. On the other hand, it has also been shown in LauS25 , that in case of cyclic sets of functional dependencies (that is if there exist $X\to A$ and $Y\to B$ in $FD$ such that $A\in Y$ and $B\in X$ hold), the set ${\sf Cons}(\mathcal{T})$ may not satisfy $FD$.
As in the case of a star-table, the set $FD$ of functional dependencies is clearly acyclic, ${\sf Cons}(\mathcal{T})$ satisfies $FD$. We thus define repairs of star-tables as follows.
Definition 2
@@
    ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$
@@
The set of all repairs of $T$ is denoted by ${\sf Rep}(T)$. $\Box$
It has been shown in LauS25 that the following holds, based on Definition 2:
@@
Elaborating briefly on the remark in the second item above, we notice that two distinct tables defined over the same universe and the same set of functional dependencies can have the same sets of true tuples. Databases such as $T$ and $R$ above are said to be equivalent and when we refer to a table $T$ we in fact refer to any table $\widetilde{T}$ such that ${\sf True}(\mathcal{T})={\sf True}(\mathcal{\widetilde{T}})$.
The following basic theorem has been shown in LauS25 to hold whenever the set $FD$ is acyclic.
@@ 4.2 Repairs of Star-Tables
Another important specificity of star-tables is that our notion of repair coincides with that in the literature, applied to ${\sf True}(\mathcal{T})$. More precisely, we show that if $T$ is a star-table and $S$ is a maximal consistent subset of ${\sf True}(\mathcal{T})$, then $S$ is a repair of $T$ in the sense of Definition 2, implying in particular that ${\sf True}(\mathcal{S})$ contains ${\sf Cons}(\mathcal{T})$. As a consequence, in Definition 2, the first item can be more simply stated as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$. This result relies on the following preliminary two lemmas.
@@
Proof. Since $S\models FD$, every $\sigma$ in the m-table $m\_Chase(S)$ is such that, for every $A$ in $sch(\sigma)$, $\sigma(A)=(a)$. Thus $m\_Chase(S)$ can be seen as a table that we denote by $S^{*}$. Moreover, denoting by $S_{t}$ the table $S\cup\{t\}$, we have $m\_Chase(S_{t})=m\_Chase(S^{*}\cup\{t\})$. Let $S^{*}_{t}=S^{*}\cup\{t\}$, and let us consider the computation of $m\_Chase(S^{*}_{t})$.
-- If $q_{1}$ and $q_{2}$ are in $S^{*}_{t}$, then either $q_{1}$ and $q_{2}$ are not defined over $A$ or, they are both defined such that $q_{1}.A=q_{2}.A$. In this case $m\_Chase$ does not change $S^{*}_{t}$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are not defined over $A$, then $m\_Chase$ does not change $S^{*}_{t}$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are defined over $A$. Since $t\in{\sf Cons}(\mathcal{T})$, it is not possible that $q_{1}.A\neq t.A$. Thus in this case again, $m\_Chase$ does not change $S^{*}_{t}$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is defined over $A$ and $q_{2}$ is not defined over $A$. Then $m\_Chase$ changes $t$ into $ta$ where $a=q_{1}.A$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is not defined over $A$ and $q_{2}$ is defined over $A$. Then $m\_Chase$ changes $q_{1}$ into $qa$ where $a=t.A$.
Indeed, let $q_{1}$ and $q_{2}$ be in $\Sigma$ and $X\to A$ be in $FD$. The only possibility for $m\_Chase$ to change $\Sigma$ is that $q_{1}.X=q_{2}.X$, $q_{1}.A=a$ and $q_{2}$ is not defined over $A$. As this does not happen in $S^{*}$ (because $m\_Chase(S^{*})=S^{*}$), and as this does not happen in $\Sigma$ (by definition of $\Sigma$), we obtain that $m\_Chase(\Sigma)=\Sigma$. It follows that $m\_Chase(S^{*}_{t})=\Sigma$, and since we have seen that no conflicts occur in $\Sigma$, $S^{*}_{t}\models FD$ holds. The proof is therefore complete. $\Box$
The following lemma is a consequence of Lemma 1.
@@
As a consequence of Proposition 3, we show that in the case of a star-table, given a true tuple $t$, there exists a repair in which this tuple is true, and that moreover, if this tuple is conflicting, there exists a repair in which this tuple is not true. It should be noticed here that, the second result is shown in LauS25 , whereas the first one is not.
@@
Proof. To show the first statement, based on the fact that for every tuple $t$ in $\mathcal{T}$, $\{t\}\models FD$ trivially holds, as ${\sf True}(\mathcal{T})$ is finite, there exists a maximal subset $S$ of ${\sf True}(\mathcal{T})$ such that $t\in S$ and $S\models FD$. By Proposition 3, $S$ is in ${\sf Rep}(T)$. Thus the proof of the first statement of the proposition is complete. The second statement being the subject of Proposition 4 of LauS25 , the proof of the proposition is complete. $\Box$

### v2

@@ 3.1 The Context of our Approach
Figure 2: The star-table $T$ and the m-table m_Chase(T)m\_Chase(T) of Example 3
In this context, the associated table $T$ over $U$ and its corresponding m-table m_Chase(T)m\_Chase(T) are shown in Figure 2. Applying Proposition 1, the sets ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$ are as follows:
  • ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of the tuples in ${\sf tuples}(\sigma)$ for every $\sigma$ in m_Chase(T)m\_Chase(T). Thus, ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of:
    - $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    - $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
    - $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
    - $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
    - $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    - $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$.
    - $k_{1}k_{2}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    - $k_{1}k^{\prime}_{2}a_{2}b_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
    - $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
    - $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$. $\Box$
3.2 m_Chasem\_Chase Computation in a Star Schema
In this section, we show that when considering a star schema, the time complexity of computing m_Chase(T)m\_Chase(T) for a star-table $T$ can be drastically lowered, as compared with the general case shown in LauS25 and recalled above. To do so, we first state important properties regarding m-tuples of m_Chase(T)m\_Chase(T) where $T$ is the table collecting all tuples in the data warehouse. In the remainder of this paper, we refer to such table as a star-table. The following proposition states important properties of the m-tuples in m_Chase(T)m\_Chase(T).
Proposition 2
Let $T$ be a star-table over universe $U$. The following hold:
  1. 1.
    For every $\sigma$ in m_Chase(T)m\_Chase(T) and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$. Consequently, if $\mathbb{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbb{K}))|=1$.
  2. 2.
    For every tuple $k$ over $\mathbb{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$.
  3. 3.
    Moreover, m_Chase(T)m\_Chase(T) contains the following two kinds of m-tuples:
    1. (a)
      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
      • $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=(k_{i_{0}})$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$,
      • for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~|~(\exists q\in D_{i_{0}})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$.
    2. (b)
      $\sigma$ such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$, and
      • for every $M_{j}\in\mathbb{M}$, $\sigma(M_{j})=\{m_{j}~|~(\exists t\in F)(t.\mathbb{K}=\sigma(\mathbb{K})\wedge t.M_{j}=m_{j})\}$,
      • for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$,
        $\sigma(A)=\{a~|~(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
Proof. See Appendix A. $\Box$
We point out that item 3 in Proposition 2 holds thanks to the restrictions on missing values earlier stated. Moreover, an important consequence of Proposition 2 is that m_Chase(T)m\_Chase(T) can be computed by an algorithm whose time complexity is significantly lower than that of Algorithm 1. This new algorithm, displayed in Algorithm 2, is shown to be correct in the following proposition.
Algorithm 2 The m-Chase Algorithm for a star-table
0: A data warehouse composed of $n$ dimensional tables $D_{1},\ldots,D_{n}$ and a fact table $F$.
0: The m-table m_Chase(T)m\_Chase(T).
1:for all $i=1,\ldots,n$ do
2:  sort the m-tuples in $D_{i}$ according to their $K_{i}$-value
3:  $D_{i}^{*}:=\emptyset$
4:  for all $K_{i}$-value $k_{i}$ occurring in $D_{i}$ do
5:   $\sigma:=(k_{i})s_{i}^{1}\ldots s_{i}^{d_{i}}$ where $s_{i}^{j}=\{a_{i}^{j}\in adom(A_{i}^{j})~|~(\exists t_{i}\in D_{i})(t_{i}.{K_{i}}=k_{i}\wedge A_{i}^{j}\in sch(t_{i})\wedge t_{i}.A_{i}^{j}=a_{i})\}$ for every $j=1,\ldots,d_{i}$
6:   $D_{i}^{*}:=D_{i}^{*}\cup\{\sigma\}$
7: sort the m-tuples in $F$ according to their $\mathbb{K}$-value
@@
9:for all $\mathbb{K}$-value $k=k_{1}\ldots k_{n}$ occurring in $F$ do
10:  $\sigma:=(k)\mu_{1}\ldots\mu_{p}$ where $\mu_{q}=\{m_{q}\in adom(M_{q})~|~(\exists t\in F)(t.{\mathbb{K}}=k\wedge M_{q}\in sch(t)\wedge t.M_{q}=m_{q})\}$ for every $q=1,\ldots,p$, and $\sigma(A)=\emptyset$ for every $A\in U\setminus(\mathbb{K}\cup\mathbb{M})$
11:  for all $i=1,\ldots,n$ do
12:   if $k_{i}$ occurs in $D_{i}^{*}$ then
13:    Let $\sigma_{i}$ be the unique m-tuple in $D_{i}^{*}$ such that $\sigma(K_{i})=(k_{i})$
14:    Mark $\sigma_{i}$ in $D_{i}^{*}$
15:    for all $j=1,\ldots,d_{i}$ do
16:     $\sigma(A_{i}^{j}):=\sigma_{i}(A_{i}^{j})$
17:  $F^{*}:=F^{*}\cup\{\sigma\}$
18:for all $i=1,\ldots,n$ do
19:  for all $\sigma_{i}$ in $D_{i}^{*}$ do
20:   if $\sigma_{i}$ is not marked then
21:    $F^{*}:=F^{*}\cup\{\sigma_{i}\}$
[…]
23:return m_Chase(T)m\_Chase(T)
Proposition 3
Given a data warehouse composed of $n$ dimensional tables $D_{1},\ldots,D_{n}$ and a fact table $F$, and its associated star-table $T$, Algorithm 2 correctly computes m_Chase(T)m\_Chase(T).
The time complexity of Algorithm 2 is in
$\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|))+|F|\cdot(\log(|F|))\right).$
Proof. It can be seen that the loop lines 18-21 in Algorithm 2 computes all m-tuples of kind (a) in Proposition 2(3), whereas the loop lines 9-17 in Algorithm 2 computes all m-tuples of kind (b). The first part of the proof is therefore complete.
Regarding the time complexity of Algorithm 2, we first notice that the statements line 2 and 7 are respectively in $\mathcal{O}(|D_{i}]\cdot\log(|D_{i}|))$ and $\mathcal{O}(|F]\cdot\log(|F|))$. Moreover, the loop lines 4-6 is linear in the size of $D_{i}$ as the table is sorted according to its key values. Thus the loop lines 1-6 is in $\mathcal{O}(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|)))$. Similarly, the loop lines 9-17 is in $\mathcal{O}(|F]\cdot\log(|F|))$ and the last loop lines 18-21 is in $\mathcal{O}(|D_{i}^{*}|)$. Since $\mathcal{O}(|D_{i}^{*}|)\approx\mathcal{O}(|D_{i}|)$, we obtain that Algorithm 2 is in $\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|))\right)+\mathcal{O}(|F]\cdot\log(|F|))+\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|)\right)$, that is in $\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|))\right)+\mathcal{O}(|F]\cdot\log(|F|))$. The proof is therefore complete. $\Box$
We emphasize that, denoting by $W$ the overall size of the underlying data warehouse (i.e., the sum of the sizes of all its tables), the complexity result in Proposition 3 can be written as $\mathcal{O}(W\cdot\log(W))$, which is significantly lower than the generic time complexity stated in LauS25 , which can be approximated to $\mathcal{O}(W^{3}\cdot\delta^{2})$. Moreover, considering Algorithm 2 instead of Algorithm 1 to assess the size of the m-table m_Chase(T)m\_Chase(T) shows that the number of m-tuples in m_Chase(T)m\_Chase(T) is exactly the number of distinct $\mathbb{K}$-values present in the fact table $F$, plus for every $i=1,\ldots,n$, the number of all $K_{i}$-values occurring in $D_{i}$ but not in $F$. On the other hand it is easy to see that, if $T$ is a star-table, the size of any m-tuple in m_Chase(T)m\_Chase(T) is bounded by $|\mathbb{K}|+\delta\cdot|U\setminus\mathbb{K}|$, which is less than $\delta\cdot|U|$ as earlier stated.
We now illustrate Proposition 3 in the context of Example 3.
$D_{1}^{*}$ $K_{1}$ $A_{1}^{1}$ $A_{1}^{2}$
$(k_{1})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$
$(k^{\prime}_{1})$ $(a_{1})$
$(k^{\prime\prime}_{1})$ $(a^{\prime}_{1})$ $(a^{\prime}_{2})$
$D_{2}^{*}$ $K_{2}$ $A_{2}^{1}$ $A_{2}^{2}$
$(k_{2})$ $(b_{1})$ $(b_{2})$
$(k^{\prime}_{2})$ $(b_{1})$
$F^{*}_{0}$ $K_{1}$ $K_{2}$ $M_{1}$
$(k_{1})$ $(k_{2})$ $(m_{1})$
$(k_{1})$ $(k^{\prime}_{2})$ $(m^{\prime}_{1}\,m^{\prime\prime}_{1})$
$(k^{\prime}_{1})$ $(k^{\prime\prime}_{2})$ $(m_{1})$
$F^{*}$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$(k_{1})$ $(k_{2})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$ $(b_{1})$ $(b_{2})$ $(m_{1})$
$(k^{\prime}_{1})$ $(k^{\prime\prime}_{2})$ $(a_{1})$ $(m_{1})$
$(k^{\prime\prime}_{1})$ $(a^{\prime}_{1})$ $(a^{\prime}_{2})$
Figure 3: The m-tables when running Algorithm 2 in the context of Example 3
@@
Applying Algorithm 2 to the tables shown in Figure 1, yields the following m-tables shown in Figure 3:
  • $D_{1}^{*}$ and $D_{2}^{*}$, when running the loop lines 1-6,
  • $F_{0}^{*}$ when considering the m-tuples $\sigma$ on line 10,
  • the first three rows in $F^{*}$ when completing the loop lines 2-5,
  • the whole m-table $F^{*}$, after computing the loop lines 18-21. Indeed, as $k^{\prime\prime}_{1}$ does not occur in $F$, the last m-tuple in $D_{1}^{*}$ is not marked because the test line 12 fails.
It is easy to see that the obtained m-table $F^{*}$ is equal to m_Chase(T)m\_Chase(T) as shown in Figure 2. Moreover, the number of m-tuples in m_Chase(T)m\_Chase(T) is $5$, corresponding respectively to the four $\mathbb{K}$-values occurring in $F$ and to the $K_{1}$-value $k_{1}^{\prime\prime}$ occurring in $D_{1}$ but not in $F$. $\Box$
@@ 4.1 Definition of Repairs
As explained in LauS25 , contrary to most approaches to consistent query answering in the literature, it is not appropriate to define a repair of $T$ as a maximal subset of ${\sf True}(\mathcal{T})$ satisfying $FD$. This is so because, although it is intuitively justified to impose that all tuples of ${\sf Cons}(\mathcal{T})$ be true in every repair, it is shown in LauS25 that in case of cyclic sets of functional dependencies (i.e., if there exist $X\to A$ and $Y\to B$ in $FD$ such that $A\in Y$ and $B\in X$), the set ${\sf Cons}(\mathcal{T})$ may not satisfy $FD$. To cope with this issue, given universe $U$, a set $FD$ of functional dependencies over $U$ and a table $T$ over $U$, a repair $R$ of $T$ is defined in LauS25 as a table $R$ over $U$ such that:
@@
    ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$.
@@
    ${\sf True}(\mathcal{R})\cap{\sf Cons}(\mathcal{T})$ is a maximal consistent subset of ${\sf Cons}(\mathcal{T})$.
  4. 4.
    For every table $R^{\prime}$ satisfying 1, 2 and 3 above, and such that ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf True}(\mathcal{R})={\sf True}(\mathcal{T})$.
Moreover, it is also shown in LauS25 that when $FD$ is acyclic, the set ${\sf Cons}(\mathcal{T})$ satisfies $FD$. Hence, it turns out from the above definition that when $FD$ is acyclic, every tuple in ${\sf Cons}(\mathcal{T})$ must be true in every repair. It is important to notice that, as shown in Example 2, even when $FD$ is acyclic, there exist maximal and consistent subsets of ${\sf True}(\mathcal{T})$ that do not contain all tuples in ${\sf Cons}(\mathcal{T})$. This shows that, even when considering acyclic sets of functional dependencies, our notion of repair does not exactly coincide with that of traditional approaches, in which repairs would be defined as maximal subsets of ${\sf True}(\mathcal{T})$ satisfying $FD$. However, we show in the next sub-section (see Proposition 4) that these two notions of repairs coincide, when considering star-tables as we do in the present approach.
It has also been shown in LauS25 that when the set $FD$ is acyclic, then the following holds:
@@
[…]
Moreover, the following basic theorem has been shown in LauS25 to hold whenever the set $FD$ is acyclic.
@@
Coming back to star-tables, the set ${\sf Cons}(\mathcal{T})$ satisfies $FD$, because as seen earlier, the set $FD$ of functional dependencies is acyclic. We thus define repairs of star-tables as follows.
Definition 2
  1. 1.
    ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$.
  2. 2.
    $R\models FD$.
  3. 3.
The set of all repairs of $T$ is denoted by ${\sf Rep}(T)$. $\Box$
It should be clear that all results recalled just above concerning acyclic sets of functional dependencies hold in the case of star-tables.
@@ 4.2 Repairs of Star-Tables
An important specificity of star-tables is that our notion of repair coincides with that in the literature, applied to ${\sf True}(\mathcal{T})$. More precisely, we show that if $T$ is a star-table and $S$ is a maximal consistent subset of ${\sf True}(\mathcal{T})$, then $S$ is a repair of $T$ in the sense of Definition 2, implying in particular that ${\sf True}(\mathcal{S})$ contains ${\sf Cons}(\mathcal{T})$. As a consequence, in Definition 2, the first item can be more simply stated as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$. This result relies on the following preliminary two lemmas.
@@
Proof. See Appendix B. $\Box$
@@
We emphasize that, Proposition 4 shows that, if ${\sf True}(\mathcal{T})$ is seen as a table over $U$, then a repair $R$ in ${\sf Rep}(T)$ can be seen as a repair of ${\sf True}(\mathcal{T})$, in the sense of traditional approaches ArenasBC99 .
On the other hand, we show below that Proposition 4 implies that, in the case of a star-table, given a true tuple $t$, there exists a repair in which $t$ is true, and that moreover, if $t$ is conflicting, there exists a repair in which $t$ is not true. It should be noticed here that, the second result is shown in LauS25 , whereas the first one is not.
@@
Proof. To show the first statement, based on the fact that for every tuple $t$ in $\mathcal{T}$, $\{t\}\models FD$ trivially holds, as ${\sf True}(\mathcal{T})$ is finite, there exists a maximal subset $S$ of ${\sf True}(\mathcal{T})$ such that $t\in S$ and $S\models FD$. By Proposition 4, $S$ is in ${\sf Rep}(T)$. Thus the proof of the first statement of the proposition is complete. The second statement of this proposition has been proved in LauS25 as Proposition 4. Therefore the proof of this proposition is complete. $\Box$